What wipl be the input A and B for boolean expression.( A+B)(A.B)=1 ..Please tell is it necessary for person who is preffering for neet has to learn boolean algebra...and questions like above....It is something which is not in text...so should i

Grade:12

1 Answers

Sanju
6 years ago
The input value is 1 for both A and B. If you put 1) A=1; B=0  2) A=0; B=1 & 3) A=0; B=0 the boolean expression will not give you 1 as output.
Explanation: When you put A=1; B=1 as input, the process goes like as follows and yield 1 as output.
Step 1: A+B = 1+1 = 1
Step 2: A.B = 1.1 = 1
Step 3:  (A+B)(A.B) = (1)(1) = 1
 
[1) INPUT A=1; B=0      OUTPUT = 0
2) INPUT A=0; B=1       OUTPUT = 0
3) INPUT A=0; B=0       OUTPUT = 0]
 
In NEET exam you can expect questions based on Boolean Algebra sinceElecronic Device portions is included in the syllabus. It is very easy to understand Boolean Algebra. Try to learn it. It will surely help you to score marks if questions are asked from that section as its very easy to crack the problems.